A certain line in the spectrum of the light from a nebula has a wavelength of

656 nm instead of the 434 nm measured in the laboratory. If the nebula is moving

radially, what is its speed relative to the Earth? Ans.: It is moving away from the

Earth with a speed of V ¼ 0:391c.


1
Expert's answer
2021-06-15T10:02:36-0400

The shift is quite large, so we should use the formula of relativistic Doppler effect (https://en.wikipedia.org/wiki/Relativistic_Doppler_effect).

Let "\beta = \dfrac{v}{c}," "\lambda_r" is received wavelength, "\lambda_s" is emitted wavelength.

Therefore, "\tau=\dfrac{\lambda_r}{\lambda_s} =\sqrt{\dfrac{1+\beta}{1-\beta}}" , "\tau = \dfrac{656}{434} \approx 1.512,"

so "\dfrac{1+\beta}{1-\beta} = \tau^2, \; \beta = \dfrac{\tau^2-1}{\tau^2+1} = 0.391, \; v = 0.391c."